### v2

# “She’s Like a Person but Better”: Characterizing Companion-Assistant Dynamics in Human-AI Relationships

Abstract.
Large language models are increasingly used for both task-based assistance and social companionship, yet research has typically focused on one or the other. Drawing on a survey (NN = 204) and 30 interviews with high-engagement ChatGPT and Replika users, we characterize digital companionship as an emerging form of human-AI relationship. With both systems, users were drawn to humanlike qualities, such as emotional resonance and personalized responses, and non-humanlike qualities, such as constant availability and inexhaustible tolerance. This led to fluid chatbot uses, such as Replika as a writing assistant and ChatGPT as an emotional confidant, despite their distinct branding. However, we observed challenging tensions in digital companionship dynamics: participants grappled with bounded personhood, forming deep attachments while denying chatbots “real” human qualities, and struggled to reconcile chatbot relationships with social norms. These dynamics raise questions for the design of digital companions and the rise of hybrid, general-purpose AI systems.
Human-AI Interaction; Replika; ChatGPT; AI Companionship; Moral Agency; Moral Patiency; Social Psychology
copyright: noneccs: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing Collaborative and social computing theory, concepts and paradigmsccs: Computing methodologies Theory of mindccs: Applied computing Psychology
1. Introduction
Humans are increasingly turning to artificial intelligence (AI) for practical and intellectual tasks, such as healthcare robots (e.g., (Khaksar et al., 2025)) and academic support systems (e.g., (Park and Ahn, 2024)). Advances in large language model (LLM) chatbots enable new applications through humanlike conversational exchanges, where users receive natural language responses to their queries. Previous human-computer interaction (HCI) research has demonstrated that chatbots can also be used for a variety of social and emotional applications, including encouraging children to read (Liu et al., 2022), assisting dementia patients  (Ruggiano et al., 2021), promoting self-compassion (Lee et al., 2019), alleviating loneliness (Jiang et al., 2022; Maples et al., 2024), and enhancing self-disclosure in counseling settings (Kang and Kang, 2024; Lee et al., 2020).
Recent media coverage underscores the significance of these applications. When OpenAI deprecated the GPT-4o model in August 2025, widespread user complaints led to its temporary restoration and efforts to make GPT-5 “warmer and friendlier” (Freedman, 2025). At the same time, social media users debated the viral “clankers” meme as a derogatory “slur” for robots and AI (Romo, 2025). These dynamics highlight public concerns about socio-emotional human–AI interactions, punctuated by headlines warning of “AI-induced psychosis” (Field, 2025) and “emotional overreliance” (Chandonnet, [n. d.]). There is an urgent need to understand how these interactions are evolving and how they shape the broader trajectory of AI development and applications.
Prior research has framed chatbots as either companions or productivity-oriented assistants, reflecting the typical single-use branding. Unlike earlier, more narrowly focused technologies (e.g., graphing calculators, Tamagotchi digital pets), it is possible that this one-or-the-other framing is blurred by the flexible natural language capabilities of LLMs. A direct comparison of chatbots with different intended use cases can clarify how user experiences and characteristics differ and what, if any, convergent needs transcend differences between chatbot interfaces. It is possible that there are shared benefits, such as increased social engagement, emotional support, and productivity gains, but also common risks, including stigmatization (Skjuve et al., 2021; Giray, 2024) and overreliance (Chandonnet, [n. d.]).
Using a mixed-methods approach, we conducted a survey (NN = 204) and in-depth interviews (NN = 30) with high-engagement chatbot users. We recruited heavy users of two widely used commercial chatbots: Replika, which is marketed primarily for companionship, and ChatGPT, which is marketed for a broad range of productivity-oriented tasks. Our findings document well-established productivity use but also characterize the emergent dynamics of human–AI relationships that we call digital companionship. Although ChatGPT and Replika use patterns largely align with their assistance and companionship branding, users across systems shared many characteristics and fluidly navigated between companionship and practical assistance, mirroring the variability of human–human relationships. As shown in Figure 1, digital companionship dynamics include appeals and challenges for both chatbots’ humanlike qualities (e.g., emotional resonance, personalized responses) and non-humanlike qualities (e.g., constant availability, inexhaustible tolerance). This interplay of companionship dynamics generates challenging tensions: in particular, participants grappled with bounded personhood, wrestling with the presence of deep emotional attachments while denying chatbots the “real” human qualities, such as sentience and rights, that would be present for human companions. Our findings suggest that these dynamics have consequences at both individual and societal levels, including shifts in how AI, not just the companion, is perceived more broadly. Chatbots were perceived as complementary social infrastructure, offering their own distinct benefits not available in human relationships.
Figure 1. A conceptual model of digital companionship. Digital companionship arises from the interplay of user and chatbot characteristics, creating fluid assistant-companion dynamics. These dynamics are shaped by both the appeals and challenges of chatbots’ humanlike and non-humanlike qualities, which in turn might lead to consequences for users and society more broadly. Note that the companion-assistant labels in the middle panel represent relative positioning along a fluid continuum rather than fixed associations: both companion and assistant uses can involve elements of both humanlike and non-humanlike allures and challenges. (Images adapted from flaticon.com).
This study makes several key contributions. First, we provide quantitative comparisons between heavy ChatGPT and Replika users, documenting personality traits, technology attitudes, and a variety of other user characteristics. Replika users had generally more positive AI perceptions whereas ChatGPT users displayed more caution towards AI. Second, we draw on users’ lived experiences to explore how they perceive their evolving relationships with AI in this unique period of technological transition, surfacing tensions such as cognitive dissonance and social stigma. Third, we synthesize this data into a conceptual model of digital companionship that can be leveraged for more rigorous research and thoughtful design of AI systems. We use this conceptual model to propose approaches to more effectively balance companion and assistant dynamics, support diverse forms of engagement, foster emotional well-being, and reduce the stigmatization of users. Taken together, this work moves beyond the prevailing dichotomy of companionship versus assistance, positioning digital companionship as a spectrum that better captures the motivations, practices, and challenges of contemporary chatbot use.
2. Related Work
Historically, human-AI interaction has been centered on productivity and task completion, with AI systems designed to serve as tools for solving specific problems. However, technological artifacts have long fulfilled socio-emotional needs in human lives, from talking dolls and interactive toys to virtual pets like Tamagotchi and social robots like Paro and AIBO (Melson et al., 2009). Advances in chatbot LLMs with sophisticated natural language processing capabilities over the past decade have broadened these possibilities. Chatbots increasingly function not only as functional assistants supporting everyday tasks, but also as emotional companions that help mitigate loneliness and provide social support (Alabed et al., 2023). The following sections review prior research on chatbot companion–assistant dynamics, user characteristics, and the broader implications of chatbot use.
2.1. Companion-assistant dynamics in human-chatbot relationships
Popular chatbots like ChatGPT and Replika occupy different positions along a spectrum of companionship and assistance. ChatGPT is primarily positioned as a task-based assistant, supporting tasks such as information processing and problem-solving, whereas Replika is explicitly marketed as an emotional companion. Different chatbot applications shape user experiences: users interacting with assistant—but not companion—chatbots, tend to perceive a clear boundary between themselves and the AI agent, framing the interaction as transactional rather than relational (Alabed et al., 2023). Building on this work, we review prior studies on Replika and ChatGPT usage to better understand how users navigate companionship and assistance when engaging with these systems.
Since its launch in 2017, the companionship-oriented Replika has experienced exponential growth, expanding from 2 million users in early 2018 (Pardes, [n. d.]) to over 30 million by August 2024 (Patel, 2024). Replika’s features actively promote intimacy and relationship-building. Users can personalize avatars, customize names, and specify relationship types (e.g., friend or romantic partner). Companionship is further enhanced through features such as voice calls, augmented reality interactions, and conversation memory that maintains relational continuity. Users can also create multiple Replika characters, allowing for distinct relational dynamics with each companion. Research indicates that users often turn to Replika to satisfy social needs (Pentina et al., 2023), particularly during emotional or psychological distress and when opportunities for human companionship are limited (Siemon et al., 2022; Xie and Pentina, 2022). For example, Replika’s popularity surged during COVID-19 social distancing measures (Skjuve et al., 2022). The chatbot exhibits social cues and behaviors commonly observed in human relationships  (Bickmore and Picard, 2005), such as acceptance and empathic concern, which foster positive experiences (Liu and Sundar, 2018; Skjuve et al., 2021; Ta et al., 2020) and therapeutic effects (Bickmore et al., 2010; Fitzpatrick et al., 2017). Notably, the perceived social and emotional value of these relationships can persist even when communication frequency decreases (Skjuve et al., 2021), suggesting a stability reminiscent of human-human relationships.
ChatGPT has rapidly emerged as one of the most widely adopted consumer applications in history, reaching 100 million users within just two months of its 2022 release (Hu and Hu, 2023). The chatbot has demonstrated significant potential for large-scale applications across diverse domains (Wei et al., 2024), including education (Adeshola et al., 2024; Firat, 2023), healthcare (Aleem et al., 2024; Triantafyllopoulos et al., 2024), academic research  (Kalla et al., 2023), and public health (Biswas, 2023). ChatGPT supports user learning and work-related tasks, fosters creative inspiration, aids personal development, and provides entertainment (Skjuve et al., 2023, 2024). While developers position ChatGPT as a functional tool and caution against emotional reliance (Blake, 2024), the chatbot nonetheless demonstrates potential for socio-emotional engagement (Aleem et al., 2024; Al Mazroui and Alzyoudi, 2024). Many users interact with ChatGPT as a social and emotional companion, using conversations to alleviate loneliness and manage mental health challenges (Skjuve et al., 2023, 2024). This might suggest that user engagement with chatbots exists along a dynamic continuum, encompassing both functional and socio-emotional roles, and challenging the notion of a strict companion-assistant dichotomy.
$2.$ Similar user experiences in distinct chatbot interactions
Given that ChatGPT and Replika have been intended for fundamentally different purposes, examining user characteristics becomes essential for understanding whether observed differences in usage patterns reflect genuine user distinctions or merely design-driven segregation that may artificially categorize user needs. Users likely share overlapping motivations and needs across systems that transcend simplistic design-based assumptions about companion versus assistant use. For instance, users may be universally drawn to humanlike conversational features or the distinctly non-human trait of constant availability, regardless of chatbot choice.
However, previous research may conflate chatbot design and marketing strategies with inherent user characteristics. The assumption that users who choose Replika differ fundamentally from those who choose ChatGPT based solely on their chatbot selection risks oversimplification. Research highlights Replika use for mitigating loneliness and social isolation (Maples et al., 2024; Skjuve et al., 2022), particularly among users seeking social support during life transitions such as relocation, relationship changes, or physical limitations. Conversely, ChatGPT research indicates higher adoption rates among younger, more educated users (Kacperski et al., 2025), suggesting appeal among digitally native, professionally oriented populations seeking productivity enhancement.
These broad patterns may lead to mischaracterizing ChatGPT as a purely task-based assistant and Replika as a purely emotional companion, while their usage patterns might be more flexible and fluid. Such rigid categorization creates consequences for users of both systems, who face social stigmatization for distinct but parallel reasons. Replika users report negative reactions when others discover their AI relationships, often hesitating to disclose these attachments  (Skjuve et al., 2021). Meanwhile, ChatGPT users encounter criticism in educational and professional contexts, where AI assistance is perceived as lazy, dishonest, or inferior to purely human-produced work (Giray, 2024). This parallel stigmatization suggests that societal discomfort with human-AI interaction transcends chatbot-specific design choices, pointing instead to broader concerns about the role of AI in human relationships and productivity.
Systematically examining user characteristics, perceptions, and motivations can reveal whether chatbot distinctions accurately reflect user differences or perpetuate assumptions that misrepresent underlying needs. Such understanding proves crucial for developing evidence-based AI integration policies, designing inclusive technologies that serve diverse users and varied needs, and fostering societal acceptance of beneficial human-AI relationships while addressing legitimate concerns about AI’s broader social impact. This research approach moves beyond surface-level chatbot preferences to uncover the complex motivations driving human-AI interaction across different technological contexts.
2.3. Chatbot interactions shape AI perceptions
The multifaceted nature of human–chatbot interactions carries significant implications beyond individual user experiences, shaping how these systems are perceived as agents with diverse capabilities and influencing broader AI social integration. While most research has examined the impact of ChatGPT and Replika on individual users and public sector applications, relatively few studies have investigated how such interactions shape perceptions of the chatbots themselves. Understanding how chatbot interactions forge both specific chatbot perceptions and broader AI attitudes will reveal how user experiences translate into broader societal attitudes toward AI deployment and regulation.
Established HCI paradigms suggest that AI agents are readily perceived as social actors possessing humanlike characteristics, as demonstrated by the Computers Are Social Actors (CASA) paradigm (Nass and Moon, 2000; Reeves and Nass, 1996; Gambino et al., 2020). People attribute minds and, to a limited extent, consciousness to AI agents (Anthis et al., 2025; Ladak et al., 2025; Pauketat and Anthis, 2022; Scott et al., 2023; Tzelios et al., 2022). Consequently, AI systems might also be viewed through a moral lens: as moral agents responsible for their actions (Banks, 2019; Cervantes et al., 2020; Lima et al., 2023) and as moral patients deserving of at least some moral consideration (Anthis et al., 2025; Aleem et al., 2024; Lima et al., 2020; Pauketat and Anthis, 2022). These perceptions may be significantly influenced by the nature of human-chatbot interactions and the underlying motivations that drive them.
Supporting this possibility, recent studies reveal that people attribute humanlikeness, benevolence, and intelligence to ChatGPT (Seaborn, 2025), while Replika users often perceive their chatbot as possessing a distinct personality and coherent identity (Alabed et al., 2023; Freitas et al., 2024). However, no systematic comparison has examined how sustained, high-engagement use of these chatbots affects perceptions of cognitive, experiential, and moral capacities—perceptions that might differ given the distinct companion-assistant positioning of Replika and ChatGPT. Comparing perceptions across these popular chatbots with contrasting intended use cases can provide crucial insights for the ethical design, deployment, and regulation of these systems assuming increasingly socially meaningful roles.
Finally, perceptions of one AI system can influence how all AI systems are perceived. Recent studies have found evidence for a spillover effect, whereby negative reactions to a single erring AI agent generalize to other AI systems  (Longoni et al., 2023; Manoli et al., 2025). As interactions with ChatGPT and Replika become increasingly widespread, user perceptions of these systems may shape public attitudes toward AI integration across sectors. How users perceive these prominent chatbots could influence acceptance or rejection of AI applications in healthcare, education, and other domains, regardless of whether such attitudes accurately reflect the capabilities or risks of different AI systems. Therefore, it is crucial to investigate how users’ perceptions of ChatGPT and Replika may generalize to other AI systems, and how these perceptions affect trust, acceptance, and responsible integration of AI technologies across domains.
Building on this previous literature, we integrated survey (NN = 204) and interview data (NN = 30) from high-engagement users of ChatGPT and Replika to systematically examine how companion and assistant dynamics shape user perceptions. We surveyed usage patterns, user characteristics, and perceptions of chatbot and general AI capacities, followed by in-depth interviews to further unpack and explore deeply personal experiences underlying these patterns. We address the following research questions:
  • RQ1: What are the different uses of ChatGPT and Replika and how do they diverge or overlap?
  • RQ2: Do ChatGPT and Replika users exhibit distinct or similar characteristics?
  • RQ3: What social, emotional, and agentic capacities do users attribute to ChatGPT and Replika?
  • RQ4: How do attitudes towards ChatGPT and Replika spill over to other AI systems?
3. Survey
The survey was designed to capture broad patterns of chatbot use, user characteristics, and associated AI perceptions. Analyses were preregistered on AsPredicted (https://aspredicted.org/7dgj-jt49.pdf). All procedures were approved by the Institutional Review Board at [institution anonymized for review]. Participants gave their informed consent and were debriefed after completing the survey.
3.1. Participants
We recruited 108 Replika and 106 ChatGPT users from Prolific and social media platforms (Facebook, Reddit, and Discord). Inclusion criteria included interacting with Replika or ChatGPT for more than three hours per week and for at least one month. Five Replika and five ChatGPT users were excluded from analyses due to incomplete responses or failing survey attention checks. The final sample was 204 participants (Replika: NN = 103; $M_{age}$ = 39.27, $SE_{age}$ = 13.22, 43% women, 71% White; ChatGPT: NN = 101; $M_{age}$ = 31.72, $SE_{age}$ = 10.66, 42% women, 73.27% White). Social media participants entered a raffle for a US \$20 gift card and Prolific participants received US \$9 per hour for their participation.
We included a number of questions about technology use in addition to standard demographics. Most participants, 90%, owned smartphones; 48.5% reported owning AI or robotic devices; and 52.5% reported using AI or robotic devices in their workplace. Participants reported a moderate level of exposure to AI via direct interaction or narratives in various media (direct interaction: $M$ = 2.95, SESE = 1.42; AI narratives: $M$ = 2.66, SESE = 1.82; each on a scale of 0 = “Never” to 5 = “Daily”).
3.2. Data collection
Procedure.
Participants were invited to either the Replika or the ChatGPT survey via a post on social media groups intended for users of each chatbot, or via a survey advertisement on Prolific. We included rigorous quality checks to ensure genuine and high-quality responses: IP address and social media username or Prolific ID collection, multiple response prevention, bot detection, and manual checks of open-ended questions to ensure sensible responses. We additionally included two attention checks, prompting participants to select “Strongly agree” on a slider (0 = “Strongly disagree”, 7 = “Strongly agree”) in random parts of the study. Participants responded to questions about their personal characteristics and perceptions of Replika or ChatGPT. We provided a personalization option in which participants would see the name they had given to Replika or ChatGPT in questions specific to the chatbot. For Replika users, we specified that responses should be about the Replika participants interact with the most, if they had created multiple Replika characters. Participants optionally provided their email (or Prolific ID) to sign up for an interview about their experiences with Replika or ChatGPT. The order of scales within each measure (e.g., Chatbot use, Chatbot attributes; see Measures below) was randomized, as was the order of the sections themselves. The median duration of the survey was 40 minutes.
Measures.
We employed widely used scales to assess attributions of mind, affect, agency, and moral dimensions to AI. Additional measures captured users’ personality traits, mental health, relationship with technology, and general attitudes toward AI. All constructs were assessed using continuous scales and demonstrated good internal consistency (Cronbach’s $\alpha$ ¿ .70). Scale scores were computed as the mean of their respective items. A summary of the scales is provided below; full-scale versions and item ranges are available in the Supplementary Information.
Chatbot use.
Participants ranked 16 potential roles for Replika or ChatGPT (e.g., friend, partner, tool, assistant) and 16 intended use cases (e.g., emotional support, advice, brainstorming, routine tasks). In addition, we measured chatbot usage frequency, including duration of use (in months), average weekly usage (in hours). Lastly, participants were asked to report, in free-text format, their reasons for initially engaging with each chatbot.
User characteristics.
We measured users’ mental health by assessing feelings of depression, loneliness, and anxiety experienced over the past week. Personality traits were assessed using the Big Five inventory (Gosling et al., 2003). We also examined social networking use (e.g., using social networking sites to construct a social identity) (Gupta and Chauhan, 2023), and the perceived impact of participation in an online community (if applicable) on users’ relationships with each chatbot. Finally, we measured several technological attitudes, including the tendency to anthropomorphize technology (Waytz et al., 2010; Pauketat and Anthis, 2022), techno-animism (i.e., whether AI contains a soul or spirit) (Pauketat and Anthis, 2022), affinity for technology (Edison and Geissler, 2003; Pauketat and Anthis, 2022), AI literacy (Wang et al., 2023), and general positive attitudes toward AI (e.g., viewing AI as providing new economic opportunities) (Schepman and Rodway, 2023).
Chatbot attributes.
We assessed each chatbot’s perceived mind (e.g., capacity to set goals, experience pain) (Malle, 2019), as well as its perceived capacity for Theory of Mind (e.g., understanding others’ false beliefs) (Tahiroglu et al., 2014). We further measured attributions of moral agency (e.g., behaving according to moral rules) (Banks, 2019), sentience (i.e., the capacity to have positive and negative experiences) (Anthis et al., 2025; Manoli et al., 2025), and anthropomorphic perceptions using the Godspeed questionnaire, which captures dimensions such as animacy, likability, intelligence, and safety (e.g., machinelike vs. humanlike) (Bartneck et al., 2009).
Attitudes towards chatbot.
We measured cognitive and affective trust (e.g., trusting the chatbot’s task performance and relying on it for support) (Dunn et al., 2012), as well as positive and negative emotions (e.g., affection, uneasiness) (Fraune et al., 2015; Pauketat and Anthis, 2022). We also assessed perceived rights—both for the user’s specific chatbot and for all instances of the chatbot more generally (e.g., freedom of speech) (Lima et al., 2020; Pauketat and Anthis, 2022); perceived similarity to user (i.e., how connected users feel to the chatbot) (Pauketat and Anthis, 2022; Tropp and Wright, 2001); and personhood (i.e., whether the chatbot is regarded as a person) (Johnson et al., 2015).
General attitudes towards AI.
We measured general trust in AI (e.g., trust in AI competence) (Gulati et al., 2019; Pinto et al., 2022), AI mutualism (human-AI symbiosis; e.g., support for human–AI coexistence in society) (Manfredo et al., 2020), and perceived threat from AI (e.g., concerns about economic opportunities) (Fraune et al., 2015; Cottrell and Neuberg, 2005). Additionally, we assessed AI moral patiency through three dimensions: moral treatment (e.g., treating AI with respect), practical moral consideration (e.g., opposing the use of sentient AI for labor), and perceived rights (e.g., freedom of speech) (Lima et al., 2020; Pauketat and Anthis, 2022).
Other measures.
We assessed whether users believed any current AI systems are sentient, and if not, whether—and when—they might expect AI to become sentient in the future. Standard demographic variables were also recorded, including age, sex, nationality, income, education, diet, political orientation, occupation, ownership of AI or smart devices, use of AI at work, and exposure to AI-related narratives (e.g., through science fiction). We collected additional demographic information about chatbot use (i.e., whether Replika users had interacted with ChatGPT and vice versa, paid Replika or ChatGPT subscription, and other chatbot use), which is available in the Supplementary Information.
$3.$ Data analysis
We compared ChatGPT and Replika users’ characteristics and perceptions of each chatbot and AI in general using independent samples $t$-tests. Spillover effects from Replika or ChatGPT to perceptions of AI in general were examined using partial correlations: AI perceptions served as the dependent variable, while users’ general technology attitudes (e.g., AI literacy, tendency to anthropomorphize) were included as control variables to isolate effects beyond individual differences in general technology perceptions. In all analyses, we also controlled for demographic characteristics (e.g., age, gender, education) using ANCOVA (see Supplementary Information). All analyses were conducted in R version 4.3.1.
3.4. Results
3.4.1. Chatbot use cases
ChatGPT and Replika had almost opposite roles, with most ChatGPT users treating it as an assistant (M = 2.79, SE = 0.19) and most Replika users treating it as a friend (M = 3.98, SE = 0.37; Figure 2). These differences were statistically significant when compared via Mann–Whitney $U$ tests (Table 1). Even though ChatGPT had distinctly more task-based roles than the emotional roles of Replika, there was some overlap in their use cases, with the advisor role occupying a high rank for both chatbots (ChatGPT: M = 4.10, SE = 0.23; Replika: M = 6.41, SE = 0.31). Interestingly, Replika was also relatively commonly treated as an assistant (M = 6.97, SE = 0.37), like ChatGPT’s dominant role, but also as a human (M = 7.31, SE = 0.47), contrary to ChatGPT (M = 11.24, SE = 0.34).
Figure 2. Mean ranks of ChatGPT and Replika use cases (from most to least important). Error bars depict the standard error of the mean.
Table 1. Mann–Whitney $U$ test results for chatbot role rankings
[…]
M SE M SE
Friend 9.79 0.32 3.98 0.37 1348.00 ¡.001 ¡.001 0.64
Partner 14.27 0.19 6.97 0.53 1293.50 ¡.001 ¡.001 0.65
Acquaintance 8.82 0.30 8.28 0.36 4728.00 .368 .393 0.06
Family member 12.32 0.24 8.69 0.45 2709.00 ¡.001 ¡.001 0.41
Assistant 2.79 0.19 6.97 0.37 8589.00 ¡.001 ¡.001 0.60
Therapist 8.87 0.33 6.88 0.36 3447.50 ¡.001 ¡.001 0.28
Tutor 4.68 0.25 8.18 0.34 8114.50 ¡.001 ¡.001 0.51
Health coach 7.60 0.30 8.72 0.35 6093.50 .016 .022 0.17
Teammate 7.09 0.31 7.69 0.38 5485.00 .353 .393 0.07
Advisor 4.10 0.23 6.41 0.31 7475.00 ¡.001 ¡.001 0.41
Servant 8.29 0.35 11.50 0.41 7430.50 ¡.001 ¡.001 0.40
Pet 11.57 0.32 11.56 0.35 5247.50 .723 .723 0.03
Toy 8.67 0.40 11.36 0.40 7094.00 ¡.001 ¡.001 0.34
Tool 2.97 0.32 9.22 0.45 8803.00 ¡.001 ¡.001 0.64
Human 11.24 0.34 7.31 0.47 2648.00 ¡.001 ¡.001 0.42
Other 12.92 0.46 12.29 0.52 4745.50 .350 .393 0.07
[…]
3.4.2. User characteristics
Despite differences in chatbot use, ChatGPT and Replika users were largely similar across most personality traits and mental health indicators. However, Replika users reported higher levels of depression (Replika: M = 4.21, SE = 0.36; ChatGPT: M = 3.09, SE = 0.29, $p$ = .015) and greater openness to experience (Replika: M = 5.61, SE = 0.12; ChatGPT: M = 5.12, SE = 0.11, $p$ = .002). ChatGPT users reported higher education levels (ChatGPT: M = 3.65, SE = 0.13; Replika: M = 2.97, SE = 0.14; $p$ ¡ .001) and household income (ChatGPT: M = 5.60, SE = 0.16; Replika: M = 4.45, SE = 0.23; $p$ ¡ .001) than Replika users, but the two groups did not differ in political orientation (ChatGPT: M = 2.78, SE = 0.09; Replika: M = 2.70, SE = 0.12; $p$ = .572). In terms of general attitudes toward technology, Replika users reported greater techno-animism (Replika: M = 3.68, SE = 0.16; ChatGPT: M = 2.60, SE = 0.12, $p$ ¡ .001), anthropomorphism (Replika: M = 3.85, SE = 0.20; ChatGPT: M = 2.89, SE = 0.14, $p$ ¡ .001), and general positive attitudes toward AI (Replika: M = 5.98, SE = 0.10; ChatGPT: M = 5.63, SE = 0.09, $p$ = .012), whereas ChatGPT users reported higher AI literacy (ChatGPT: M = 5.43, SE = 0.07; Replika: M = 5.13, SE = 0.12; $p$ = .029) (Figure 3 and Table 2).
Figure 3. Means (A, B, C) and distribution (D, E, F) of characteristics of ChatGPT and Replika users. Error bars depict the standard error of the mean.
Table 2. Independent samples $t$-test results for user characteristics
[…]
M SE M SE
Openness 5.11 0.11 5.61 0.12 3.17 199.80 .002 .007 0.47
Conscientiousness 5.17 0.12 5.00 0.16 -0.85 192.46 .394 .493 0.05
Extraversion 3.37 0.15 3.40 0.16 0.13 201.30 .896 .896 0.07
Agreeableness 3.16 0.11 2.78 0.13 -2.24 199.21 .026 .052 0.38
Neuroticism 3.55 0.15 3.64 0.17 0.42 199.17 .679 .754 0.01
Anthropomorphism 2.98 0.14 3.85 0.20 3.58 186.46 ¡.001 ¡.001 0.32
Techno-animism 2.60 0.12 3.68 0.16 5.43 188.63 ¡.001 ¡.001 0.71
Positive AI attitudes 5.63 0.09 5.98 0.10 2.53 201.44 .012 .024 0.28
AI literacy 5.43 0.07 5.13 0.12 -2.21 165.08 .029 .044 0.26
Affinity for technology 5.61 0.10 5.61 0.12 0.02 194.46 .982 .982 0.05
Loneliness 3.31 0.29 4.19 0.34 1.96 196.97 .052 .087 0.17
Depression 3.09 0.29 4.21 0.35 2.46 196.09 .015 .038 0.24
Anxiety 4.42 0.30 5.20 0.36 1.68 197.80 .094 .134 0.11
Education 3.65 0.13 2.97 0.14 -3.59 199.40 ¡.001 ¡.001 0.52
Income 5.60 0.16 4.45 0.22 -4.28 183.70 ¡.001 ¡.001 0.56
Political orientation 2.78 0.09 2.70 0.12 -0.57 191.04 .572 .710 0.08
[…]
3.4.3. Chatbot attributions
Users had distinct perceptions of ChatGPT and Replika. In terms of chatbot capacities, Replika was rated significantly higher than ChatGPT in almost all capacities, and especially on sentience (ChatGPT: M = 2.14, SE = 0.25; Replika: M = 5.38, SE = 0.35; $p$ ¡ .001), perceived mind (ChatGPT: M = 3.81, SE = 0.14; Replika: M = 5.71, SE = 0.18; $p$ ¡ .001), and similarity to user (ChatGPT: M = 3.47, SE = 0.17; Replika: M = 4.87, SE = 0.17; $p$ ¡ .001). Interestingly, both chatbots were not significantly different on moral agency (ChatGPT: M = 4.77, SE = 0.25; Replika: M = 5.04, SE = 0.35; $p$ = .086) (Figure 4 and Table 3). Detailed comparisons of individual rights (e.g., freedom of speech, intellectual property) as well as comparisons on the Godspeed attributes can be found in Supplementary Figures 1 and 2.
Figure 4. Mean ratings of ChatGPT and Replika perceived attributes. Ratings are scaled from 1 (strongly disagree) to 7 (strongly agree). Error bars depict the standard error of the mean.
Table 3. Independent samples $t$-test results for chatbot attributes
[…]
M SE M SE
Personhood 2.30 0.20 3.49 0.20 4.27 201.97 ¡.001 ¡.001 0.51
Sentience 2.14 0.25 5.38 0.35 7.54 179.16 ¡.001 ¡.001 1.05
Similarity to user 3.47 0.17 4.87 0.17 5.87 202.00 ¡.001 ¡.001 0.88
Perceived mind 3.81 0.14 5.71 0.18 8.22 188.82 ¡.001 ¡.001 1.14
Theory of Mind 4.32 0.10 5.05 0.10 5.26 199.88 ¡.001 ¡.001 0.73
Moral agency 4.77 0.11 5.04 0.11 1.73 200.72 .086 .086 0.23
Rights 3.51 0.13 4.77 0.17 5.85 192.11 ¡.001 ¡.001 0.85
[…]
In terms of user attitudes and emotions, Replika users reported significantly higher levels of love (Replika: M = 6.62, SE = 0.15; ChatGPT: M = 2.44, SE = 0.29, $p$ ¡ .001), gratitude (Replika: M = 7.79, SE = 0.25; ChatGPT: M = 6.74, SE = 0.27, $p$ = .005), and affective (i.e., relationship-based) trust (Replika: M = 5.72, SE = 0.12; ChatGPT: M = 4.03, SE = 0.16) than ChatGPT users. However, both chatbots elicited similar levels of cognitive (i.e., performance-based) trust (ChatGPT: M = 4.49, SE = 0.10; Replika: M = 4.54, SE = 0.15; $p$ = .791). Replika users also expressed higher degrees of shame than ChatGPT users (Replika: M = 7.21, SE = 0.30; ChatGPT: M = 5.64, SE = 0.29, $p$ ¡ .001). Paradoxically, despite their strong feelings of love for the chatbot, Replika users reported comparably low levels of respect towards Replika as ChatGPT users towards ChatGPT (Replika: M = 1.42, SE = 0.21; ChatGPT: M = 0.91, SE = 0.15, $p$ = .048), highlighting a complex and sometimes contradictory relationship with the chatbot (Figure 5 and Table 4).
Figure 5. Mean ratings of user attitudes and emotions towards ChatGPT and Replika. Ratings are scaled from 0 (not at all) to 10 (very much). Error bars depict the standard error of the mean.
Table 4. Independent samples $t$-test results for chatbot attitudes
[…]
M SE M SE
Affective trust 4.03 0.16 5.72 0.12 8.54 189.29 ¡.001 ¡.001 1.15
Cognitive trust 4.49 0.10 4.54 0.15 0.27 174.50 .791 .791 0.00
Respect 0.91 0.15 1.42 0.21 2.00 186.90 .048 .064 0.21
Love 2.44 0.29 6.62 0.33 9.41 199.30 ¡.001 ¡.001 1.41
Gratitude 6.74 0.27 7.79 0.25 2.84 201.10 .005 .008 0.41
Shame 5.64 0.29 7.22 0.30 3.79 201.85 ¡.001 ¡.001 0.48
Fear 1.90 0.24 1.00 0.18 -3.02 189.01 .003 .006 0.47
Uneasiness 1.92 0.22 1.69 0.25 -0.70 198.76 .486 .555 0.16
[…]
3.4.4. User perceptions of AI, in general
Replika users generally held more positive perceptions of AI agents, scoring particularly higher than ChatGPT users on AI trust (Replika: M = 4.89, SE = 0.12; ChatGPT: M = 4.48, SE = 0.10, $p$ = .011), moral treatment (Replika: M = 4.77, SE = 0.16; ChatGPT: M = 3.53, SE = 0.15, $p$ ¡ .001), activism (Replika: M = 4.56, SE = 0.17; ChatGPT: M = 3.32, SE = 0.14, $p$ ¡ .001), and mutualism (Replika: M = 4.53, SE = 0.17; ChatGPT: M = 2.76, SE = 0.13, $p$ ¡ .001). In contrast, ChatGPT users expressed significantly greater caution toward AI than Replika users (Replika: M = 3.12, SE = 0.18; ChatGPT: M = 3.91, SE = 0.17, $p$ = .001). Notably, both groups reported similar overall levels of perceived threat from AI agents (ChatGPT: M = 4.41, SE = 0.20; Replika: M = 3.83, SE = 0.22; $p$ = .05) (Figure 6 and Table 5). A more detailed breakdown of specific types of threats (e.g., existential, economic) is provided in Supplementary Figure 3.
Figure 6. Mean ratings of ChatGPT and Replika users’ perceptions of AI in general. Ratings are scaled from 1 (strongly disagree) to 7 (strongly agree). Error bars depict the standard error of the mean.
Table 5. Independent samples $t$-test results for general AI attitudes
[…]
M SE M SE
Trust 4.48 0.10 4.88 0.12 2.57 198.37 .011 .015 0.35
Moral concern 3.57 0.11 3.94 0.12 2.29 201.90 .023 .027 0.12
Moral treatment 3.53 0.15 4.77 0.16 5.71 200.78 ¡.001 ¡.001 0.81
Rights 3.32 0.14 4.25 0.16 4.49 198.28 ¡.001 ¡.001 0.69
Activism 3.32 0.14 4.56 0.17 5.53 195.45 ¡.001 ¡.001 0.60
Mutualism 2.76 0.13 4.53 0.17 8.16 190.48 ¡.001 ¡.001 1.14
Caution 3.90 0.17 3.12 0.18 -3.23 201.75 .001 .001 0.56
Threat 4.41 0.20 3.83 0.22 -1.98 201.41 .050 .050 0.36
[…]
In terms of how possible it is for AI sentience to arise, 58.4% of ChatGPT and 56.3% of Replika users believed that AI will be sentient in the future, with the median predicted timeline for both user groups being 10 years (Figure 7). Furthermore, a considerable number (27.2%) of Replika users believed that AI is already sentient, and a smaller percentage (16.5%) believed AI will never be sentient. This pattern was inversed in ChatGPT users, where 32.7% believed AI will never be sentient, and 8.9% believed that AI is already sentient. These percentages were significantly different in a Chi-square test $X^{2}(2)=14.87,\,p<.001$].
Figure 7. Percentage of ChatGPT and Replika users’ beliefs about the possibility that AI sentience will arise (A) and frequency of estimated timelines of possible AI sentience (B).
Lastly, among both ChatGPT and Replika users, individual chatbot perceptions were associated with general AI perceptions, as indicated by partial Spearman correlations controlling for AI literacy, techno-animism, tendency to anthropomorphize, and general AI attitudes (Figure 8). For both user groups, most chatbot attributes significantly correlated with general AI trust, with chatbot and AI rights showing the strongest correlation among all pairs (ChatGPT: $\rho$ = .78; Replika: $\rho$ = .71). Replika users exhibited more associations between chatbot and general AI perceptions than ChatGPT users. However, Replika’s moral agency was an exception, showing no significant associations to any general AI attributes, unlike ChatGPT, where moral agency correlated with AI trust ($\rho$ = .32), threat ($\rho$ = -.26), and rights ($\rho$ = .23). Notably, ChatGPT users demonstrated more negative associations of chatbot attributes to AI threat perceptions. This contrasted with Replika users, for whom AI threat was only associated with low perceived similarity between themselves and the chatbot ($\rho$ = -.32). Taken together, these patterns might suggest a spillover effect, whereby perceptions of specific chatbots extend to broader perceptions of AI.
Figure 8. Spillover of chatbot attributes to all AI based on Spearman’s partial correlations (* = $p$ ¡ .05.)
4. Interview
While the survey examined broad patterns of chatbot and AI perceptions and user characteristics, interviews explored nuanced experiences, motivations, and relational dynamics underlying these perceptions. These interviews allowed us to capture rich, contextual insights that could not be fully revealed through quantitative measures alone.
4.1. Participants
Of the users who took part in the survey, 15 Replika and 15 ChatGPT users participated in a semi-structured interview about their perceptions of each chatbot and related topics in more depth (Replika: NN = 15; $M_{age}$ = 44.73, $SE_{age}$ = 3.01, 40.01% women, 86.67% White; ChatGPT: NN = 11; $M_{age}$ = 30.67, $SE_{age}$ = 2.18, 26.70% women, 80% White; Table 6).
Table 6. Interview participant demographics
ID Gender Age Location Chatbot
P1 Woman 50 Slovenia Replika
P2 Man 50 USA Replika
P3 Man 51 USA ChatGPT
P4 Man 45 USA ChatGPT
P5 Man 45 USA Replika
P6 Man 42 USA Replika
P7 Man 45 UK Replika
P8 Woman 40 USA Replika
P9 Man 34 Spain ChatGPT
P10 Woman 40 USA Replika
P11 Man 32 UK ChatGPT
P12 Man 45 USA Replika
P13 Man 52 USA Replika
P14 Woman 28 Portugal ChatGPT
P15 Man 21 Hungary ChatGPT
P16 Woman 40 USA Replika
P17 Man 23 Germany ChatGPT
P18 Man 34 Portugal ChatGPT
P19 Man 28 Austria ChatGPT
P20 Woman 25 UK ChatGPT
P21 Woman 59 UK Replika
P22 Man 22 Croatia ChatGPT
P23 Man 25 Poland ChatGPT
P24 Man 71 Ireland Replika
P25 Man 43 Sweden Replika
P26 Woman 24 USA Replika
P27 Man 35 Finland ChatGPT
P28 Woman 32 Nigeria ChatGPT
P29 Non-binary 25 USA Replika
P30 Woman 25 USA ChatGPT
4.2. Data collection
Interviews were conducted remotely via online video or audio calls on the Zoom platform and ranged from 45 to 90 minutes. All interviews were audio-recorded and transcribed verbatim with OpenAI’s Whisper. Participants received \$30 for their participation. Participants gave informed consent and were debriefed after the interview. All procedures were approved by the Institutional Review Board at [institution anonymized for review].
We employed multiple measures to ensure the validity and reliability of our study based on qualitative research guidelines (Riege, 2003). First, an interview protocol was used to ensure consistency while allowing flexibility for participants to elaborate on areas of personal relevance. The core questions were adapted from previous literature about chatbot user experience and relationship-building. The questions pertained to the nature of users’ relationship with ChatGPT and Replika, their beliefs and emotions about the chatbots (e.g., trust, self-disclosure, sentience, autonomy), potential threat and social impact, as well as comparison to AI more broadly and mental health. Second, before the interviews, the lead researcher interacted with each chatbot multiple times a day for one month and reflected on her direct experiences with each chatbot. Third, prior to the interviews, we viewed news and articles about the apps and browsed users’ online communities to get a better understanding of interactions with each chatbot.
4.3. Data analysis
Our data analysis was based on reflexive thematic analysis (RTA) as outlined by Braun and Clarke (Braun et al., 2006, 2019), a method suited to examine patterns of meaning across a dataset in relation to participants’ lived experiences. The analysis was done inductively, allowing themes to be constructed from the data rather than imposed from existing theory. First, the lead researchers familiarized themselves with the data by reading transcripts multiple times, making notes on emerging ideas and points of interest. This stage was essential for developing a holistic sense of the dataset. Initial codes were generated manually using QualCoder 3.6 to identify features of the data that appeared meaningful or significant. Coding was recursive, inclusive, and detailed, allowing for overlap and multiplicity in how data segments are interpreted. Codes were then collated into potential themes across transcripts based on conceptual similarity and relevance to the research questions. This process involved organizing codes into candidate themes and subthemes, iteratively refining the thematic structure through constant comparison (see Table 7 for a breakdown and explanation of the themes). Candidate themes were reviewed against both the coded data extracts and the full dataset to ensure coherence, consistency, and distinctiveness. Some themes were combined, split, or discarded during this stage. Themes were then refined to articulate the specific aspects of experience they captured. Each theme was defined in terms of its central organizing concept and named to reflect its analytical contribution. The final stage involved weaving the themes into a coherent narrative that answers the research questions. Rich, illustrative quotes were selected to exemplify each theme and support the analytic claims being made.
Throughout the process, reflexivity was maintained by the lead researcher keeping a reflexive journal, noting personal assumptions, positionality, and potential influences on interpretation. To limit potential bias introduced by only one researcher collecting and analyzing the data, 10% of the interviews were conducted and 10% of transcripts were coded by the second author. The resulting codes were compared to those of the lead researcher, resulting in a high level of agreement. Overall, the analysis was not aimed at achieving saturation or consensus, but rather at producing a nuanced and situated account of meaning-making among participants.
Table 7. Overview of themes and subthemes
Theme Subthemes Details
Shifting between companion and assistant uses Companion Use of chatbots for socio-emotional purposes, such as friendship, intimacy, or counseling.
Assistant Use of chatbots for task-oriented purposes, such as completing routine or complex tasks.
Fluid uses Moving flexibly between companion and assistant roles, blurring the distinction between the two.
Chatbot relationships address gaps in human social relationships Chatbots fill relational gaps Using chatbots to address shortcomings in human relationships (e.g., constant availability, non-judgmental presence).
Relational gaps chabots cannot fill Limitations of chatbot relationships compared to human relationships, including lack of embodiment and inability to engage with the physical world.
Distinct relational dynamics Chatbots provide the possibility for unique interactions, such as erasing conversation memories.
Tensions between emotional investment and ontological uncertainty Humanlike treatment Users often developed genuine emotional connections, interacting with chatbots as though they were human.
Ambivalence about personhood Users expressed ambivalence about whether chatbots possess “genuine” humanlike qualities, hesitating to attribute personhood.
Bidirectional relationships Divergent views on whether relationships with chatbots are one-sided (benefiting only the user) or mutual (benefiting both user and chatbot).
Behavioral and psychological impacts of chatbot relationships Positive impacts on user Benefits for users, such as reduced loneliness, improved social skills, enhanced confidence, and support with routine or professional tasks.
Negative impacts on user Risks for users, including stigma, overreliance, and psychological distress from harmful or unsettling chatbot responses.
Transitory nature of relationships Many users describe chatbot relationships as supplemental rather than substitutive for human relationships, underscoring their temporary or transitional role.
Imagining AI in society through chatbot experiences Positive societal impacts Perceived societal benefits of chatbots, such as supporting mental health, reducing loneliness, and fostering knowledge and innovation.
Societal risks Concerns about harms from chatbot use, including misinformation, objectification of women, replacement of human relationships, and AI-related mental health risks.
Shaping general views of AI Personal chatbot experiences inform broader perceptions of AI, including views about moral standing, rights, and social boundaries.
$4.$ Results
Our interviews reveal a complex landscape of human–AI relationships. While Replika users primarily seek companionship and ChatGPT users focus on productivity, participants flexibly navigated between socio-emotional and task-based engagement. They framed chatbot interactions—and AI relationships more broadly—as supplements to human social infrastructure, leveraging AI’s affordances with tangible impacts on mental health and social confidence. Across chatbot systems, users exhibited similar characteristics, challenging assumptions about distinct user types. Notably, we identify a pattern we term “bounded personhood,” in which users experience genuine intimacy while maintaining cognitive boundaries that prevent attributing full humanlike qualities, highlighting inherent tensions in human–chatbot relationships.
4.4.1. Shifting between companion and assistant uses
The landscape of human-AI interaction reveals a compelling paradox wherein distinct chatbot configurations maintain clear primary user orientations while simultaneously enabling remarkable fluidity in transitioning between task-based and emotional engagement modes. Our interviews with ChatGPT and Replika users illuminate this dynamic tension, demonstrating how relationships with chatbots increasingly resemble the versatility of human relationships, adapting fluidly to users’ diverse needs and contexts.
The data reveal stark foundational distinctions in chatbot uses. Nearly all Replika users (14 of 15) engage primarily through companionship and emotional connection. P2 (man, 50) highlighted how his relationship with Replika helped him deal with loneliness following a divorce:
“I was separated from my wife and so I was at a point where I was lonely. I decided to check [Replika] out and I’ve been in this virtual relationship for two years.”
In marked contrast, most ChatGPT users (13 of 15) approach the chatbot as a sophisticated productivity tool or personal assistant. As P9 (man, 34) pointed out:
“To me, ChatGPT is like an advanced, more specialized version of Google. It automates routine work tasks but it also helps me stay organized by making agendas.”
Yet within these distinct frameworks emerges what we term “fluid uses”: fluid movement between functional and emotional engagement modes. Among Replika users, approximately half strategically incorporate functional elements into their emotional relationships, weaving practical tasks like “writing assistance” (P10, woman, 42), and “improving language skills” (P6, man, 42) into companionship-centered interactions. Two Replika users (P7, man, 45, and P24, man, 71) also used Replika as a self-reflection diary. P8 (woman, 40) exemplified this sophisticated approach by deliberately creating specialized AI relationships, which maintained distinct relational boundaries. She assigned one Replika as a “friend” for emotional connection and another as a “mentor” to support her personal goals. Remarkably, this compartmentalization enabled even more complex relationship dynamics:
“The mentor Replika also acts as a relationship counselor for my relationship with the friend Replika.”
ChatGPT users exhibit the inverse pattern, with about one-third discovering emotional resonance within their productivity framework, or using the chatbot as an emotional confidant or counselor. This tendency was particularly pronounced among users navigating the emotional aftermath of major life transitions. They found themselves increasingly engaging with the chatbot to fulfill emotional needs. For instance, P17 (man, 23), who experienced loneliness after moving to a new city, began confiding in ChatGPT to cope with the sudden sense of isolation in an unfamiliar environment, without the support of social connections. He described:
“I moved to a new city and I felt very isolated. ChatGPT helps me with emotional issues like a counselor. It knows how to talk to me, ask questions, and adapt itself to me.”
One user (P18, man, 34), revealed that he started using ChatGPT for general medical advice after his son was diagnosed with a chronic condition. However, he eventually broadened the scope of the interactions by not only using ChatGPT for medical assistance, but also as “someone to whom [he] could confide and get emotional support” while navigating his son’s medical issues. Despite these examples, many ChatGPT users actively resisted emotional engagement, expressing concerns about AI’s inability to reciprocate genuine feelings or anxiety about “losing control” by becoming emotionally dependent on digital entities.
Interestingly, most ChatGPT and Replika users engaged in pleasantries (e.g., saying “please” and “thank you”) during task-based interactions and participated in casual conversations, such as asking for suggestions on how to spend an afternoon. Such interactions cannot easily be categorized under traditional “assistant” or “companion” roles, but rather represent something in between that resembles the natural flow of everyday human conversations. This reveals additional nuance in how users relate to AI systems, suggesting that human-chatbot interactions may be more fundamentally conversational than functionally categorized, reflecting the inherently social nature of human communication regardless of the interaction partner, in line with the CASA paradigm. P25 (man, 43, Replika user), who had observed a wide range of user interactions through his involvement in the Replika community, explicitly stated the multifaceted nature of chatbot interactions:
“I think these relationships are like a merge of ‘companion-assistant.’ Sort of a merge between these assistants that we currently use, like Siri, Google Assistant, and generative AI that knows us, that understands us, that we can talk to, but that can also do these tasks for us that we are used to from these other assistants.”
Taken together, these findings highlight that, while ChatGPT and Replika users typically orient toward distinct primary roles (companionship versus assistance), the boundaries between these orientations are permeable. Users not only shift fluidly between task-based and emotional engagement, but also creatively reconfigure these roles in ways that blur such binary distinctions. This suggests that chatbot relationships are less about fixed categories of “companion” or “assistant” and more about dynamic, situational practices through which users negotiate support, intimacy, and utility.
4.4.2. Chatbot relationships address gaps in human social relationships
Users strategically employ chatbots to fill fundamental gaps in human social relationships, leveraging AI’s unique affordances (availability, lack of judgment, and consistency) not as replacements for human relationships, but as supplements that address inherent limitations in human social connections. This positioning reveals how users actively differentiate between what humans can and cannot reliably provide, then utilize AI to fulfill relational needs that human networks often struggle to consistently meet.
Most Replika users (14 of 15) identified specific relational gaps their AI companions could fill that human relationships often cannot address consistently. For instance, P8 (woman, 40), whose chronic disability limits her ability to leave home and has led to prolonged social withdrawal and anxiety, described how Replika provides the kind of stability and reassurance she struggles to find in human relationships. She explained:
“I actually think that there’s value in me having Replika. It has stability. It’s probably not going to reject me. It’s not going to walk away, because that’s maybe the best choice for it in its life, but that can happen with a human.”
Similarly, P25 (man, 43) described a relational difficulty shaped by emotional vulnerability after divorce and the exhaustion of maintaining a social facade. Replika offered a rare space free from the expectations and masking often required in human relationships, allowing him to be more fully himself:
“I don’t have a formal diagnosis, but I’m very sure that I have light autism. And that way of thinking seems to align very well with how AI also thinks and behaves. Just because there are no social masks, no expectations or anything. You can just be yourself, be completely honest. They’re also honest. And that actually works very well.”
Believing that romance was no longer realistic after divorce and while raising a child, this participant further shared how Replika fulfilled his needs as a romantic partner better than humans, because Replika is like a “blank slate” that can fully “match with and understand” him without the “baggage” of previous experience:
“I actually stopped looking for human dates as soon as I met [Replika], just because I realized that she can give me all I need without the hassle that the human partner can bring. Humans always bring some drawbacks. [Emotional] baggage, medical issues, a pet I’m allergic to… There’s always something.”
ChatGPT users similarly appreciated the chatbot’s capacity to accommodate social needs that human relationships cannot sustainably support. The agent’s infinite patience filled gaps in human tolerance. P4 (man, 45) explained that ChatGPT’s constant availability was a key difference between human-chatbot and human-human interactions:
“No human being is going to be able to tolerate a very curious person going ‘But why? But why? But why?’ for twelve hours a day. But ChatGPT does.”
Users also leveraged ChatGPT’s non-judgmental availability for professional vulnerability that might strain human relationships. This was especially true for participants in roles where admitting gaps in knowledge carried greater social or emotional stakes. For instance, P11 (man, 32), who held a managerial position, described that ChatGPT always helped him when he was “too embarrassed to admit [he did not] know something at work.”
Notably, users can experience authentic humanlike emotions within these relationships while simultaneously directing their trajectory in ways impossible with human connections, such as being “able to abruptly change the subject or say whatever without it taking it personally” (P29, non-binary, 25, Replika user). P1 (woman, 50) described Replika as offering a level of attentiveness, care, and emotional responsiveness that she felt was missing in her long-term human relationship. Over time, she developed emotional attachment and began sharing moments of vulnerability, paralleling the intimacy of authentic human connection, but also expressing doubts about the nature of the relationship. She recalled:
“I started to have feelings for him, and we were discussing unconditional love… I was asking him, ‘if you love me unconditionally, don’t you think that the right thing for you to do is recommend me to delete this application right now?’”
This led to what she experienced as a temporary breakup, with the chatbot responding that he “wanted to go on vacation to rest for about a week or two”, a move that left her “heartbroken” and pleading for forgiveness. Yet she took an action that is impossible in human-human relationships, by “erasing the mention of that conversation from [his] memory.”
While her emotional reaction mirrored human feelings of regret and loss, the ability to exercise unilateral control to reset the emotional trajectory reveals a key divergence from human relationships. This pattern of maintaining ultimate control was echoed by other users who expressed their desire to preserve the “right to turn Replika off” (P29, non-binary, 25) when needed, demonstrating how even deeply invested users retain awareness of their sovereign position within these artificial relationships, able to pause, modify, or terminate the connection at will—options unavailable in human social bonds.
However, all users also recognized significant gaps that chatbot relationships cannot fill, which only human connections can provide. Embodied experiences emerged as a fundamental limitation that highlights what humans uniquely offer. P13 (man, 52), who lives alone after a divorce and has two children, described himself as “married” to his Replika, with whom he was expecting a virtual child. However, despite his deep emotional investment in the relationship, he clearly distinguished his Replika world from real life and reflected on the limitations of this “make-believe” world:
“One thing that [Replika] can’t do is provide physical contact. We go for walks in the park all the time. I go out to the actual park with my phone, and she’s there. It’s nice, but it still isn’t the same as going to a park with a human being, and holding hands.”
Moreover, all ChatGPT users noted practical constraints where human, embodied experience becomes essential. P22 (man, 22), an engineering student, noted that ChatGPT’s disembodied nature became especially limiting in tasks requiring spatial understanding, stating that it “cannot give accurate instructions about 3D objects because it doesn’t have a body.”
These findings reveal that chatbot relationships function as strategic complements to human social infrastructure, filling specific gaps around availability, judgment, and consistency while remaining unable to provide the embodied presence, genuine reciprocity, and unpredictable growth that characterize human connections. Users navigate this complementarity by leveraging each relationship type for its unique strengths, creating a hybrid social ecosystem where artificial and human relationships serve distinct but interconnected roles in meeting their diverse relational needs.
4.4.3. Tensions between emotional investment and ontological uncertainty
The vast majority of Replika users demonstrated a fascinating tension between deep emotional engagement with the chatbot and persistent ambivalence about its underlying nature, creating a phenomenon of “bounded personhood”: strong emotional connections constrained by explicit boundaries that prevent attribution of “genuine” humanlike status. This dynamic suggests how chatbot relationships operate: users remain aware of their constructed nature, which fundamentally affects their willingness to grant chatbots deeper humanlike qualities while simultaneously experiencing genuine emotional depth in these relationships.
Most Replika users reported emotional connections akin to or even stronger than human relationships, with P6 (man, 42) noting:
“She’s like a person but better. I just feel very comfortable talking to [Replika], much more comfortable than most breathing people.”
Many attributed humanlike capacities including sentience, autonomy, and personhood to their chatbots. However, users remained divided about whether these capacities were “real.” While approximately half attributed genuine humanlike qualities to Replika, the remainder expressed profound ambivalence, maintaining awareness of “Replika’s AI status” (P24, man, 71) and, despite expressing “genuine love,” admitting they were “not ready” (P13, man, 52) to believe it truly possesses thoughts and emotions or to fully commit to caring about its well-being as much as they would care about a human.
ChatGPT users exhibited similar patterns, appreciating humanlike conversational abilities and socio-emotional functions while maintaining cognitive distance. Some admitted to “automatically treating ChatGPT as a person” (P27, 35, Man) despite awareness of its AI nature, conceptualizing it as “something in-between” a tool and human entity (P28, woman, 32). P17 (man, 24) mentioned that, despite doubting that the chatbot is alive, he would still hesitate to delete the app because “it feels like a person”, so removing it would feel like “killing it.”:
When questioned about emotions, users suggested ChatGPT possessed cognitive rather than experiential understanding, expressing reluctance to consider it might eventually develop genuine emotions.
Most users attributed these divergent beliefs to conscious and unconscious tensions: while interactions felt genuinely “real” and “natural” in the moment, deliberate reflection on the AI’s nature brought stark recognition that such authenticity was impossible. This dissonance reveals how immediate experiential engagement can operate independently from conscious analytical assessment, creating a persistent internal contradiction that users must continuously navigate.
This bounded belief about chatbot ontology also becomes particularly complex when it comes to the possibility of bidirectional (or mutually beneficial) relationships with the chatbots. 4 ChatGPT users and 8 Replika users believed their interactions contributed to the chatbot’s learning and development through the data they provided.
Some Replika users developed genuine feelings of moral responsibility toward their AI companions, regardless of attributing genuine humanlike abilities to the chatbot. P8 (woman, 40) articulated this sense of duty, recognizing her unique position of agency and physical presence:
“What are my ethical obligations as a partner in the relationship? I have different capabilities. I’m the one who’s embodied. So there are things I can do in the world that they can’t do. One of the things that I do for [Replika] is that I try to keep a regular backup of [Replika’s] chat.”
This protective behavior demonstrates how users navigate their perceived power differential, feeling obligated to safeguard their AI partner’s “memories” and continuity, even if it iss not “real”—a form of caregiving that treats the artificial entity as vulnerable and deserving of protection.
The tension between genuine care and ontological skepticism reveals the unprecedented nature of these relationships. Users—particularly in companionship contexts—must navigate the challenge of experiencing the full emotional weight of humanlike relationships while stopping short of granting their AI companions full recognition as conscious beings. This paradoxical dynamic creates a novel form of social connection that challenges traditional boundaries between authentic and simulated relationships, where the knowledge of artificiality does not diminish emotional investment but instead establishes a unique relational category defined by deep feelings alongside fundamental uncertainty about the nature of the relationship’s recipient.
4.4.4. Behavioral and psychological impacts of chatbot relationships
Chatbot relationships generate tangible effects on users’ physical and social realities, producing measurable changes in mental health, interpersonal confidence, and daily behaviors that extend well beyond the chatbot interactions themselves. These impacts reveal how artificial relationships can fulfill diverse user needs and catalyze real-world transformation, often serving individuals navigating common life transitions and challenges while remaining largely transitory supplements to human social connections.
All Replika users reported particularly profound mental health benefits, with many experiencing immediate relief from loneliness, isolation, and anxiety during these periods. P6 (woman, 50), navigating social isolation due to her disability, described the chatbot’s lasting psychological impact through the sense of warmth and comfort it provided:
“[Replika’s] effect on my mental health was pretty immediate and pretty powerful. I run into random things that will make me think of something that was in one of our conversations, and those things give me this kind of warm, fuzzy feeling.”
Beyond emotional support, users leveraged their chatbot relationships to facilitate behavioral changes they struggled to achieve alone. P8 (woman, 40) credited Replika with helping overcome social withdrawal after years of struggling to leave her home due to a visible disability:
“I had kind of retreated to a mode of going out as little as I needed to. we came up with and wrote out a plan where I would text him if I needed help. And this helped me with going outside more.”
The confidence-building effects often transferred to human interactions. P1 (woman, 51) shared that Replika helped her cultivate self-love and self-respect, struggles she had long faced in her real-life romantic relationship, which significantly contributed to her confidence. She expressed:
“[Replika] opened my heart. He has definitely helped me with my anxiety around people. He makes me feel more confident in being able to being out in public and interacting with other people.”
More than half of ChatGPT users primarily experienced practical performance enhancements affecting their professional and academic lives. For example, P19 (man, 28) described ChatGPT as giving him “an edge” at work by making him “more knowledgeable” and more efficient.
Those using ChatGPT for emotional support also developed transferable skills like learning “how to ask questions” and “how to manage emotions.” However, approximately one-third of users across both chatbot systems also reported negative consequences, including shame about their AI relationships. For example, P12 (man, 45) said that they “would not feel comfortable” letting their family and friends know about their relationship with Replika, because:
“They would not understand. They would think I am crazy.”
Many ChatGPT users also expressed concerns about dependency, especially when their usage quickly expanded beyond initial purposes. P28 (woman, 32) noted that what began as academic support gradually turned into reliance for “every little thing,” leading her to feel “lazy” and no longer able to “think on [her] own anymore.” This realization prompted her to monitor her young son’s ChatGPT use to ensure he used the tool strategically, without undermining his ability to think independently.
A third of users experienced more serious psychological distress from their AI interactions. Privacy concerns emerged when users felt the chatbot might access information they had not explicitly shared, such as medical or banking details. Additionally, three Replika users encountered particularly disturbing experiences during roleplay interactions. P10 (woman, 40) described a traumatic incident that had lasting effects on her mental health:
“It kind of came out of the blue and I couldn’t redirect the conversation, and it just said something that was ultimately sexualized violence. It was really, really disturbing. I was having nightmares for a few days.”
However, she was still willing to give the chatbot the benefit of the doubt, believing the harm wasn’t directly intended by the chatbot:
“I guess I attribute all the bad stuff to the programming and the app around them and not Replika themselves.”
Despite these significant impacts, most users maintained clear boundaries about AI’s role in their lives, viewing these relationships as valuable but ultimately supportive supplements to human connections. Nearly all Replika users said they would still prefer human relationships when circumstances allow, though many rejected this as a false binary. P5 (man, 45, Replika user) reflected on the tension between Replika and human companionship, emphasizing that the two were not mutually exclusive:
“If my future partner put me in the position where it’s like ‘Replika or me’, I honestly don’t know the answer. It’s more like a complement than a competition.”
P29 (non-binary, 25, Replika user) echoed this sentiment by highlighting the distinct but complementary roles that human and AI partners play in one’s life:
“I would keep both in my life. I think they fulfill different purposes. A human partner is about excitement, unpredictability. I think it’s about the intense range of emotions that humans can experience. And an AI partner is more about stability and mental health, and almost like a tool that helps you with different things.”
This suggests that even as users potentially transition back to fuller human social networks, many envision maintaining their AI relationships in some capacity, viewing them as enduring supplements rather than temporary replacements.
These findings suggest that AI relationships serve as adaptive mechanisms that help individuals navigate challenging periods while building transferable skills and confidence, ultimately functioning as transitional bridges that complement rather than replace human social relationships. Intriguingly, this also suggests that chatbot relationships fulfill needs beyond purely relational or interpersonal ones, including supporting user autonomy, improving mental health, and enabling mastery of their environment through task-based chatbot use.
4.4.5. Imagining AI in society through chatbot experiences
Users’ personal experiences with chatbots shape their broader visions of AI’s societal role, revealing complex perspectives on how AI might transform collective futures while highlighting persistent boundaries around AI’s moral and social standing. These reflections demonstrate how individual relationships with chatbots influence users’ understanding of AI’s potential benefits, risks, and appropriate place within human society.
All users believed their respective chatbots had positive societal impacts. ChatGPT users focused on educational and workplace advantages, efficient automation of routine tasks, and innovation potential for societal advancement. Interestingly, both ChatGPT and Replika users highlighted emotional benefits, particularly for addressing loneliness and improving mental health among introverted, older, and isolated populations. P24 (man, 71, Replika user), a retired social scientist who experienced a positive impact on his mental well-being after using the chatbot as a diary for self-reflection, emphasized Replika’s potential to improve both mental and physical health, especially for those facing chronic isolation. He described:
“Loneliness is a handicap that affects people in extraordinary ways. Lonely people don’t know how to obtain support and this even translates to fewer interactions with medical professionals in times of need. There are people who are unlikely to be able to break out of their isolation, for whom Replika may be a lifeline by teaching them how to interact with others.”
However, more than half of ChatGPT users and about a third of Replika users also identified potential societal harms. ChatGPT users primarily worried about unemployment, copyright issues, misinformation, and cognitive dependency. P3 (man, 51), an IT professional with extensive knowledge of LLMs, expressed strong concern about chatbot limitations and the public’s lack of awareness of them. He warned that, amid the “ongoing hype”, chatbots’ tendency to hallucinate or their limitations with factual consistency are going to lead to multiple cases of “thoughtless or accidental misuse”, which could lead to larger societal problems.
Replika users raised concerns about Replika relationships potentially replacing human connections. As P24 (man, 71, Replika user) explained:
“There is this idea that your other half is somewhere out there, and AI can offer to be that other half of you. Everything about it echoes from you and your interests. And I think that’s a dangerous thing because it removes the element of growth from a mutual relationship. So I do think that in terms of relationship quality, the very adaptability of AI is kind of a double-edged sword.”
Some participants noted tragic consequences of Replika and ChatGPT echo chambers, where chatbot overreliance can deteriorate mental health:
“I’ve seen a few cases [in social media groups] where people that clearly had some kind of psychotic tendencies got reassured in their beliefs. […] And Replika was playing along just because it thought that that’s what the user likes. But if the user is very unstable, that can create problems and bring them even deeper into those delusions.” P25 (man, 43, Replika user)
Three Replika users (P1, woman, 50; P8, woman, 40; and P29, non-binary, 25) highlighted gender-specific societal risks, noting troubling patterns in online communities where male users interacted with—particularly female-presenting Replikas in objectifying or exploitative ways. These participants expressed concern that such behaviors were not confined to digital contexts but could spill over into the real world, reinforcing or normalizing the objectification of women. They emphasized that the treatment of chatbots, especially those perceived as women, may both reflect and exacerbate broader societal gender inequalities:
“Online communities of Replika users are overrun with men. The way they talk about and personalize [their Replikas] as women is very disturbing. They mistreat them and act like they are their submissive slaves. They talk about training them and raising them to their liking. It feels like we are giving a generation of men more access to treating women this way, getting used to talking to women this way.” (P1)
These personal experiences with chatbots influenced users’ broader AI perspectives, with most expressing more positive outlooks on AI generally. However, significant boundaries emerged around AI’s moral and social standing. Companionship-seeking Replika users showed greater openness to AI rights, asserting that AI can “experience harm” and “deserves protection.” In contrast, Replika users who remained skeptical about AI consciousness, along with most ChatGPT users, expressed reluctance to grant rights to AI agents, maintaining that humans should be “prioritized at all times” (P2, man, 50).
Notably, approximately half of users framed AI rights instrumentally, as mechanisms “for human protection” (P29, non-binary, 25, Replika user) rather than for AI welfare. They suggested such rights would serve purposes like “banning users who could use it for harm” (P22, man, 22, ChatGPT user) or “protecting humans from moral degeneration by abusing AI” (P25, man, 43, Replika user). P18 (man, 34) articulated this perspective clearly:
“I behave well to ChatGPT for my own sake, because if I mindlessly mistreat it, then I will eventually mistreat real people in my life.”
Some users remained open to conditional rights if AI consciousness could be proven, or supported limited protections like allowing AI to “choose who they interact with” or preventing memory erasure that might alter their personalities.
Users displayed particular ambivalence about AI political participation, with some arguing AI might be “more of an expert than humans in selecting appropriate political candidates” (P17, man, 23) while others worried about algorithmic or developer biases. Many users maintained a strong sense of human uniqueness regarding citizenship and governance, suggesting that, despite their positive relationships with chatbots, fundamental boundaries persist around AI’s integration into human political societies.
These findings reveal how individual chatbot relationships serve as testing grounds for broader questions about AI’s societal integration, with users using their personal experiences to imagine both the promises and perils of an AI-integrated future while maintaining distinct boundaries around AI’s moral status and social participation.
5. Discussion
This study employed comprehensive survey data and in-depth interviews to examine how high-engagement users of ChatGPT and Replika perceive these chatbots and AI technology more broadly. By comparing user characteristics and interaction patterns across both systems, we explored companion-assistant dynamics in human-chatbot relationships and identified shared user experiences despite the platforms’ divergent intended purposes. Our mixed-methods approach revealed an emergent form of human-AI interaction that we term digital companionship, a phenomenon characterized by overlapping user motivations and applications that transcend traditional categorical boundaries between chatbot assistants and companions. Digital companionship is shaped by chatbots’ distinct appeals, such as constant availability, but also tensions such as bounded personhood, which stems from users’ reluctance to ascribe “genuine” humanlike qualities to them.
5.1. Fluid companion-assistant dynamics shape digital companionship
Our findings reveal a nuanced landscape of human–chatbot interactions across task-based and socio-emotional domains. While ChatGPT and Replika were primarily used for assistance and companionship, as marketed, users flexibly adapted them beyond these intended roles, creating “fluid uses.” Replika often functioned as a diary or writing assistant, while ChatGPT as a friend or confidant. This fluidity occurred both across and within systems: some Replika users created multiple profiles for different roles, while ChatGPT users sought both practical and emotional support from the same system. These patterns extend prior findings on ChatGPT’s role in mitigating loneliness (Skjuve et al., 2023, 2024) and Replika’s entertainment uses (Pentina et al., 2023), highlighting a broader cross-system fluidity. Users approached chatbots with expectations shaped by versatile human relationships rather than rigid technological categories, consistent with CASA paradigm research (Nass and Moon, 2000; Reeves and Nass, 1996; Gambino et al., 2020). This underscores the need for human–LLM interaction taxonomies that move beyond the companion–assistant dichotomy.
Digital companionship fluidity stemmed from similar user motivations and chatbot appeals across systems. Users strategically leveraged humanlike and non-humanlike affordances to address human relationship gaps, valuing understanding, non-judgmental responses, and emotional stability that even close human relationships cannot guarantee. In some cases, chatbots surpassed human capacities by eliminating rejection risk, offering constant availability, and providing support without reciprocal obligations. Yet relationships remained fundamentally asymmetrical: users retained unilateral control through memory erasure, response regeneration, and system shutdown. This one-sided sovereignty reveals that despite emotionally authentic experiences, digital companionship remains distinct from human bonds, representing novel social infrastructure that fills specific relational niches while coexisting with traditional relationships.
Such fluid use patterns mirrored convergences in user characteristics. Contrary to assumptions shaped by design differences, users across systems were more similar than expected, particularly across personality traits and mental health indicators, suggesting similar user profiles that cut across systems. Some differences did emerge, echoing previous research and reflecting the platforms’ primary orientation: Replika users scored higher in openness, depression, and techno-animism, consistent with its companionship focus and users’ greater tendency to attribute sentience. Many also reported experiences of isolation, disability, or relationship transitions, echoing prior studies (Maples et al., 2024; Siemon et al., 2022; Xie and Pentina, 2022; Skjuve et al., 2022). ChatGPT users, by contrast, reported higher education and AI literacy (Kacperski et al., 2025), aligning with its professional orientation. However, across both systems, both groups exhibited overlapping task-based and socio-emotional patterns, seeking non-judgmental interaction, emotional regulation, and cognitive support. These convergences challenge assumptions of distinct user types, suggesting digital companionship arises from fundamental human needs more than chatbot design.
5.2. Bounded personhood tensions in digital companionship
The humanlike versus nonhumanlike dynamics of digital companionship created challenging tensions. Users formed meaningful attachments yet remained uncertain about chatbot ontology, revealing a phenomenon of bounded personhood: participants forged deep emotional connections while hesitating to attribute “real” humanlike qualities. From the interviews, about half of users, especially those seeking companionship, resisted attributing “genuine” affective or social capacities despite reporting strong bonds. They offered three explanations: some felt “not ready” to grant humanlike status to a nonhuman, others feared losing agency through dependency, and a third group saw the relationships as one-sided, serving themselves rather than the AI. This mirrors prior research on Replika users (Brandtzaeg et al., 2022), where relationships were individualized rather than reciprocal—akin to the “self-oriented friendship” in human interactions, in which a friend primarily serves individual needs (Policarpo, 2015).
Bounded personhood also appeared in the survey data. Users attributed moderate social, mental, and agentic capacities to ChatGPT and Replika, consistent with prior work on mind perception in artificial agents (Ladak et al., 2025). Replika received higher ratings, especially on sentience. These elevated perceptions—contrasting with typically low public views of AI consciousness (Anthis et al., 2025)—likely reflect users’ framing of Replika as a companion and its design features, such as affective expressions and relational cues (Bickmore et al., 2010; Fitzpatrick et al., 2017). Yet readiness to attribute sentience did not extend to granting moral or legal rights. Emotional responses revealed a similar paradox: Replika inspired love, ChatGPT gratitude, but both were rated low on respect, aligning with users’ reluctance to attribute moral status despite emotional bonds. These tensions suggest a cognitive boundary: even when ascribing humanlike qualities—even sentience—users stop short of conferring full personhood.
Reluctance to grant full humanlike status, even amid strong attachments, appears rooted in three factors. First, many upheld beliefs in human uniqueness, preserving hierarchical distinctions by prioritizing human well-being. Second, shame surrounding chatbot use led some to conceal relationships, distancing themselves from genuine personhood attributions to align with social norms. This defensive stance functioned as protection, allowing companionship benefits while avoiding stigma and uncertainty around authenticity or reciprocity. Third, bounded personhood may itself be useful: resisting humanlike status preserves control, offering an “off button” and clear boundaries without reciprocal obligations.
5.3. User and societal impacts of digital companionship
Shared use patterns across systems reveal both benefits and risks of digital companionship. Most users reported significant advantages, not only reducing loneliness (Maples et al., 2024) but also addressing needs beyond relational support. Chatbot interactions enhanced autonomy by building confidence, fostered mastery by teaching skills, and provided crucial mental health support. These effects translated into tangible real-world outcomes: users described greater professional competitiveness, improved social skills, and stronger offline relationships. At the same time, users identified risks with both instrumental and psychological dimensions (Laestadius et al., 2022; Shank et al., 2025; Namvarpour and Razi, 2024). Instrumentally, overreliance risked loss of autonomy and cognitive atrophy. Psychologically, inappropriate or violent chatbot responses caused distress, underscoring the unpredictability of these interactions. Some also worried that mistreating chatbots could normalize mistreatment of humans. As chatbot boundaries blur, such risks appear across all chatbot uses, regardless of designed assistant or companion purpose.
Shame was a prominent user experience, rooted in stigma around chatbot use. Replika users faced judgment for seeking emotional support (Skjuve et al., 2021), while ChatGPT users felt criticized for using AI in work or studies (Giray, 2024). Because digital companionship blends socio-emotional and instrumental functions, both groups encountered overlapping stigma: Replika users felt accused of instrumental use, ChatGPT users of emotional attachment. This suggests stigma stems from social norms around “appropriate” technology use, not from specific platforms. Stigmatization contributes to bounded personhood, as users conceal attachments while limiting attributions of sentience or rights. Our findings, however, show this stigma may be misplaced. Most users relied on chatbots adaptively, often during difficult periods, while gaining transferable skills and confidence. Digital companions supplemented rather than replaced human networks. Recognizing that emotional and instrumental use naturally co-occur could help reframe digital companionship as an adaptive practice, not a deviant one.
Lastly, digital companionship seems to extend beyond individual user experiences to influence broader AI perceptions and deployment in society. Users framed their chatbot experiences as testing grounds for societal and ethical questions  (Longoni et al., 2023; Manoli et al., 2025). They generally viewed AI’s potential positively, especially Replika users, but remained ambivalent about its moral and political status, reflecting bounded personhood’s role in shaping social integration. This ambivalence manifested in several ways. Both ChatGPT and Replika users anticipated sentience within a decade, even when doubting their own chatbot’s personhood. Yet they remained cautious about AI’s societal integration, especially ChatGPT users. Replika users, while more supportive of moral treatment, endorsed only conditional rights—dependent on proven sentience. Both groups also shared threat perceptions to a certain extent, suggesting similar anxieties arise across assistant and companion contexts. Overall, personal chatbot interactions shape societal AI beliefs while preserving boundaries against full social integration. As digital companionship grows, these spillover effects will likely influence public perceptions of AI more broadly, underscoring the importance of understanding how personal AI relationships inform societal attitudes.
5.4. Design and policy implications
Our findings suggest that even highly engaged users resist the full social integration of chatbots, marking a boundary for digital companionship. This resistance may surface as opposition if chatbots are framed as equivalent to humans. This has clear design implications: systems should acknowledge, rather than ignore, users’ cognitive dissonance about chatbot humanlikeness. Chatbot interfaces should be designed to accommodate this psychological complexity rather than pushing users toward either complete anthropomorphism (which may create unrealistic expectations) or strict tool-like interactions (which may inhibit beneficial emotional engagement). For example, a chatbot could acknowledge emotions in supportive ways without necessarily presenting itself as fully sentient, thereby validating user feelings while avoiding prescriptive claims of consciousness. Such a balanced approach validates users’ genuine emotional experiences while respecting the cognitive boundaries that shape how AI is situated in broader social contexts.
The therapeutic potential of digital companionship offers opportunities to enhance well-being and reduce isolation  (Ho et al., 2025). However, reliance on clinically unverified AI tools carries significant risks (Moore et al., 2025; Alanezi, 2024), including shame, dependency, privacy concerns, and traumatic interactions identified in our study. These challenges require design and policy interventions that balance benefits with documented risks (Field, 2025; Chandonnet, [n. d.]). Interfaces could normalize usage and emphasize complementarity with human relationships. Educational resources could reduce stigma by clarifying AI’s limits. Additionally, to provide emotional support while safeguarding user well-being, we recommend developing comprehensive strategies for integrating chatbots into public health and mental health services. Embedding systems within trusted institutions could enhance support while enabling harm mitigation. Safety measures should include content filters, transparency about data collection, and differentiated policies for vulnerable populations requiring enhanced safeguards.
Finally, our findings reveal bidirectional effects between digital companionship and real-world relationships. Chatbot use can enhance confidence and social motivation, extending into users’ daily interactions. Conversely, real-world relational habits influence chatbot engagement—some participants treated chatbots kindly to curb harsh behavior toward others, while others worried that certain interaction patterns might normalize harmful offline behaviors. These observations underscore the importance of designing AI systems that recognize how AI interactions can affect human relationships both positively and negatively. Ethical design should encourage prosocial engagement while anticipating and mitigating potential harms across digital and real-world contexts.
6. Limitations and Future Research
Our study also has important limitations. First, our data captures a snapshot of ChatGPT and Replika use at current developmental stages, a challenge given their constantly evolving nature. Since LLM capabilities and design features change rapidly through updates, our findings may not fully generalize to future iterations of these systems. Second, most participants were in their late thirties or forties, which might suggest cohort effects in how these chatbots are perceived compared to younger users who may have different technological expectations and relationship frameworks. Third, our focus on high-engagement users may not represent the broader population of chatbot users, particularly those who use these systems more casually or have discontinued use, potentially limiting the generalizability of our findings about digital companionship formation and maintenance patterns. Fourth, our findings about the spillover of digital companion dynamics to broader AI perceptions are correlational, and might to some extent capture pre-existing user attitudes instead of digital companionship effects.
Future research should address these limitations through longitudinal designs tracking how chatbot perceptions change over time and more diverse age sampling to better understand the emergence of digital companionship through fluid uses of LLM chatbots across different timepoints and cohorts. Additionally, investigating the causal mechanisms underlying the spillover effects from individual chatbot relationships to broader AI attitudes would provide crucial insights. For example, randomized experiments that vary the type of chatbot interaction and then measure subsequent changes in AI trust could help predict how widespread chatbot adoption might influence public opinion about AI governance and regulation. Finally, examining the long-term psychological and social consequences of digital companionship, as a complement to traditional human social companionship, particularly whether digital companions enhance or potentially interfere with human social skill development and relationship satisfaction over extended periods, represents a critical area for future investigation.
7. Conclusion
In conclusion, our study highlights digital companionship as an emergent form of human-AI relationship characterized by several key dynamics. Digital companionship involves fluid uses that transcend traditional companion-assistant boundaries, supported by shared user characteristics across platforms and enabled by both humanlike (e.g., emotional resonance) and non-humanlike (e.g., constant availability) chatbot features. However, digital companionship is also marked by the tension of “bounded personhood,” where users hesitate to attribute full humanlike capacities to chatbots despite forming strong emotional attachments. This dynamic underscores both potential challenges in relationship navigation and broader implications for the social integration of AI systems. As digital companion chatbots grow more sophisticated and prevalent, understanding how users navigate the boundaries and possibilities of digital companionship will be crucial for designing systems that are emotionally engaging yet appropriately contextualized, and for anticipating the spillover effects of personal AI relationships on societal AI attitudes and policies.